OpenStudy (anonymous):

Linear algebra question: Find the minimum angle between the vectors: 3i+j and -i+j How do I have to reason?

myininaya (myininaya):

Do you know how to find the dot product of the two vectors given?

myininaya (myininaya):

you will have to use the formula \[\theta=\cos^{-1}(\frac{u \cdot v}{||u|| \cdot ||v||})\]

OpenStudy (anonymous):

a·b = ||a|| ||b|| cos(theta)

OpenStudy (anonymous):

All right...

OpenStudy (anonymous):

If I do the calculations I get \[\cos^{-1} (\frac{ -1 }{ \sqrt{5} })\]

OpenStudy (anonymous):

Can you check if it's correct? Thank you for your help.

myininaya (myininaya):

looks good

OpenStudy (anonymous):

ah ok. Thank you very much :)